Use the law of exponents to simplify.
6^3/4 multiply 6^1/8
would it be 6^3/32?
Any help would be appreciated.

Answer by richwmiller(17219)

No
let's take a simple example
2^6=64
2^5=32
2^4=16
2^3=8
2^2=4
8*4=32
2^3*2^2=32
which is 2^5
so we add exponents when we multiply numbers with the same base

(6^(3/4))*( 6^(1/8)=6^(3/4+1/8)=6^(6/8+1/8)=6^(7/8)
(6^(3/4))^1/8=6^(3/32) where we are raising the power to a power